French Language Definition File for master bibliographic style files

Patrick W. Daly

This paper describes file french.mbs, version 1.1 from 1995/05/04

1 Introduction

This file is to be used as a language definition file with master bibliographic files. The language it defines is

French

This is not a self-sufficient .mbs file, but rather may be used together with one that contains language support via external definition files.

See the documentation for makebst for formatting details.

2 The Menu Information

All the menu information is nested between docstrip guard options %<*options> … %</options>, and the last command is \endoptions. The rest of the file

is nested between %<*!options&!driver> . . . %</!options&!driver> in order

to exclude it if docstrip is used to extract only the menu information or the documentation driver.

1 ⟨∗options⟩

 2 \mes{^^JOptions for FRENCH}
 3 \wr{\spsp\spsp\sep french\pc: French language}

4

 5 \mes{^^JABBREVIATE WORD `PAGES':}
 6 \optdef{*}{}{`Page(s)'}{(no abbreviation)}
 7 \optdef{a}{pp}{`Page' abbreviated}{as p. or pp.}
 8 \getans

9

10 \mes{^^JABBREVIATE WORD `REDAKTEUR':}
11 \optdef{*}{}{`R\string\'edakteur(s)'}{(no abbreviation)}
12 \optdef{a}{ed}{`R\string\'edakteur' abbreviated}{as R\string\'ed.}
13 \getans

14

15 \mes{^^JOTHER ABBREVIATIONS:}
16 \optdef{*}{}{No abbreviations}{of Tome, \string\'edition, chapitre}

2 P. W. Daly

17 \optdef{a}{abr}{Abbreviations}{of such words}
18 \getans

19

20 \mes{^^JEDITION NUMBERS:}
21 \optdef{*}{}{Write out editions}{as primi\string\`ere, deuxi\string\`eme,...}
22 \optdef{n}{ord}{Numerical editions}{as 1re, 2e, 3e, etc}
23 \getans

24

25 \endoptions

26 ⟨/options⟩